Project

General

Profile

Watchdog Notes 11/18:

Attending: Dom, Nick, Michael, Kirk, Mat, Kanika, Gavin, Siva

Operations:
Power outage over the weekend due to false water leak, cold diblock 3.

Production: We seem to have NUmI files missing at Fermilab starting around 11690. (Looks like this was due to beam being off.)
Nick- Bad Channels: How do we use it? MInerba’s scheme still works. (We have just one mask.)
-The cosmics masks exist - to apply we need a hack.
: Jon’s bad channel system is not working…. Jon’s working on this. (Gav talked to Jon last week.)
: Gavin is going to summarize the issues from Friday and email Jon.

Michael: Automating NUMI scanning process. Gavin, Jim and Mark will open this up in the next day or two. Watchdogs could start scanning soon.
- Watchdog Wednesday is tutorial for scanning.

Kanika: Speak to Steve about shift form:
- Some run have no information.
- Bring it up at 3 pm.
- What the status of Joao’s work?

Key tab issues: This will prevent nearline plots until it's resolved.